6. A 0.09 m 3 drum contains saturated liquid-vapor mixture at 280 O C. Find the mass of each

phase if their volumes are equal and what is the percent quality of the mixture.


1
Expert's answer
2021-05-05T07:40:38-0400

at "T = 280^0C"

"V_f = 0.001332 \frac {m^3}{kg}" and "V_{fg} =0.03017 \frac{m^3}{kg}" From Steam table

"V =0.09 m^3"



"x =\frac{V_f} {V_fg}"


"x =\frac{0.03017} {0.09} =0.3352 = 33.52" %